OpenStudy (anonymous):

Write a polynomial function of least degree with integral coefficients that has the given zeros. -2, -8, -7

OpenStudy (anonymous):

A zero is a solution to a given function when f(x) = 0. First, let's factor in our zeros. \[f(x) = 0 = \left( x + 2 \right)\left( x + 7 \right)\left( x + 8 \right)\] \[f(x) = (x ^{2} + 9x + 14)\left( x + 8 \right)\] \[f(x) = x ^{3} + 17x ^{2} + 86x + 112\] And there you have it.

OpenStudy (anonymous):

Further clarification: in this case, zeros are used in this format \[-2 = x \rightarrow \left( x + 2 \right) = 0\] \[-8 = x \rightarrow \left( x + 8 \right) = 0\] \[-7 = x \rightarrow \left( x + 7 \right) = 0\] is it easier to see how, 'zeros' are basically solutions to a function when it equates to zero?